two fishing boats depart a harbor at the same time, one traveling east, the other south. the eastbound boat travels at a speed 3 mi/h faster then the southbound boat. after two hours the boats are 30 mi apart. find the speed of the southbound boat
Answer by mananth(16946)

they are travelling at right angles to each other.
At any given instant they form a right triangle with their starting point

speed of DISABLED_event_one= x South bound
other speed = x+3 East bound
after two hours
they will be 2x & 2(x+3) miles away from the starting point
(2x)^2+(2(x+3))^2=900
4x^2+(2x+6)^2=900
4x^2+4x^2+24x+36=900
8x^2+24x-864=0
/8
x^2+3x-108=0
x^2+12x-9x-108=0
x(x+12)-9(x+12)=0
(x+12)(x-9)=0
take the postive value
x= 9 mph the speed of south bound
12mph east bound